### v1

# Predicting random close packing of binary hard-disk mixtures via third-virial-based parameters

Abstract
We propose a simple and accurate approach to estimate the random close packing (RCP) fraction of binary hard-disk mixtures. By introducing a parameter based on the reduced third virial coefficient of the mixture, we show that the RCP fraction depends nearly linearly on this parameter, leading to a universal collapse of simulation data across a wide range of size ratios and compositions. Comparisons with previous models by Brouwers and Zaccone demonstrate that our approach provides the most consistent and accurate predictions. The method can be naturally extended to polydisperse mixtures with continuous size distributions, offering a robust framework for understanding the universality of RCP in hard-disk systems.
I Introduction
Ever since the introduction of the concept by Bernal [1, 2, 3, 4, 5], random close packing (RCP) has remained a subject of sustained interest across mathematics, physics, and engineering [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. This enduring attention stems from the ubiquitous appearance of densely packed disordered structures in a wide variety of systems, including jamming transitions, colloids, granular media, glasses, amorphous solids, ceramics, and even living matter, as illustrated by the extensive body of work cited above.
A commonly adopted definition of RCP—although sometimes challenged as ill-defined—identifies it as the highest achievable density at which a large collection of particles can be packed without long-range order. The concept was originally formulated for monodisperse hard spheres and later extended to monodisperse hard disks, systems that have played a paradigmatic role in the study of packing problems. An additional complication arises from the fact that different packing protocols, whether numerical or experimental, may lead to different maximal densities [27, 37].
Polydispersity further affects the RCP fraction, which helps explain why studies of polydisperse HS and HD systems are comparatively scarce relative to their monodisperse counterparts. From a theoretical standpoint, the analytic determination of the RCP fraction in such systems remains essentially an open problem. As a result, recent theoretical efforts have focused on approximate, physically motivated descriptions aimed at capturing the dominant effects of size disparity in simple mixtures.
In two recent papers [63, 64], Brouwers proposed, using a geometric approach, the following approximation for the RCP fraction ($\phi_{\text{mixt}}$) of a binary HD mixture with small or moderate dispersity:
$\phi_{\text{mixt}}=\frac{\phi_{\text{mono}}}{1-\mu_{B}(1-\phi_{\text{mono}})},\quad\mu_{B}\equiv\frac{(q^{2}-1)^{2}}{2(q^{2}+1)}\frac{x_{\text{L}}x_{\text{S}}}{m_{2}},$ (1a)
$\frac{\phi_{\text{mixt}}}{1-\phi_{\text{mixt}}}=\lambda_{B}\frac{\phi_{\text{mono}}}{1-\phi_{\text{mono}}},\quad\lambda_{B}\equiv\frac{1}{1-\mu_{B}}.$ (1b)
Here, $\phi_{\text{mono}}$ denotes the RCP fraction of the monodisperse HD system, $x_{\text{L}}$ and $x_{\text{S}}=1-x_{\text{L}}$ are the mole fractions of large and small disks, respectively, $q\equiv\sigma_{\text{L}}/\sigma_{\text{S}}\geq 1$ is the size ratio, and $m_{n}=M_{n}/M_{1}^{n}$ are reduced moments, with $M_{n}=x_{\text{S}}\sigma_{\text{S}}^{n}+x_{\text{L}}\sigma_{\text{L}}^{n}$ the moments of the size distribution.
By fitting Eq. (1a) to Monte Carlo simulation data [73] for $q=1.4$, Brouwers obtained $\phi_{\text{mono}}=0.8425$, a value slightly below the commonly accepted estimate $\phi_{\text{mono}}=0.844$ [37].
Figure 1: (a) $1/\phi_{\text{mixt}}$ versus the parameter $\mu_{B}$ [Eq. (1a)] and (b) $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ versus $\lambda_{B}$ [Eq. (1b)] for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data [73], while the dashed and dashed-dotted lines correspond to Eqs. (1a) and (1b) with $\phi_{\text{mono}}=0.844$ and $\phi_{\text{mono}}=0.8425$, respectively.
Two noteworthy consequences follow from Eqs. (1). First, the entire dependence of $\phi_{\text{mixt}}$ on $x_{\text{L}}$ and $q$ is encoded in the single parameter $\mu_{B}$ (or, equivalently, in $\lambda_{B}$). Consequently, two distinct binary mixtures sharing the same value of $\mu_{B}$ are predicted to have the same RCP fraction. Second, the inverse RCP fraction $1/\phi_{\text{mixt}}$ depends linearly on $\mu_{B}$, while the ratio $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ depends linearly on $\lambda_{B}$. As shown in the Appendix, these two linear dependencies uniquely fix the functional form of Eqs. (1). This implies that plots of $1/\phi_{\text{mixt}}$ versus $\mu_{B}$ and $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ versus $\lambda_{B}$ for different mixtures should exhibit data collapse.
However, as illustrated in Fig. 1, the collapse of simulation data corresponding to different mixtures is imperfect, particularly as the size ratio $q$ increases.
The main purpose of this short paper is to demonstrate that a significantly better collapse is achieved by slightly modifying Eqs. (1) and introducing an alternative “universality” parameter based on the third virial coefficient of the mixture. We further demonstrate that this approach can be naturally extended to polydisperse mixtures, providing a universal and practically useful framework for predicting RCP fractions in both discrete and continuous size distributions.
Figure 2: (a) $\phi_{\text{mixt}}$ versus the parameter $\mu$ [Eq. (2a)] and (b) $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ versus $\lambda$ [Eq. (2b)] for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data [73], and the dashed lines correspond to our proposal [Eqs. (2)] with $\phi_{\text{mono}}=0.844$.
Figure 3: $\phi_{\text{mixt}}$ as a function of the area fraction of large disks, $c_{\text{L}}$, for binary HD mixtures with size ratios $q=1.4$, $1.7$, 22, and $3$. Symbols are simulation data [73]. Solid lines show our predictions [Eq. (2a)] with $\phi_{\text{mono}}=0.844$, while the dashed-dotted and dashed lines correspond to Eq. (1a) with $\phi_{\text{mono}}=0.8425$ and Eq. (II) with $\phi_{\text{mono}}=0.844$, respectively.
II Our proposal
In our approach, we replace Eqs. (1) with the following expressions:
$\phi_{\text{mixt}}={\phi_{\text{mono}}}+\mu(1-\phi_{\text{mono}}),\quad\mu\equiv\frac{b_{3}-1-\left(\bar{B}_{3}-1\right)m_{2}}{b_{3}-3},$ (2a)
$\frac{\phi_{\text{mixt}}}{1-\phi_{\text{mixt}}}=\frac{\lambda}{1-\phi_{\text{mono}}}-1,\quad\lambda\equiv\frac{1}{1-\mu},$ (2b)
where $b_{3}=4\left(\frac{4}{3}-\frac{\sqrt{3}}{\pi}\right)\simeq 3.12802$, and
$\bar{B}_{3}=$ $\frac{1}{m_{2}^{2}}\left[x_{\text{S}}^{3}b_{3}+\frac{x_{\text{S}}^{2}x_{\text{L}}}{\pi}F(q)+\frac{x_{\text{S}}x_{\text{L}}^{2}}{\pi}q^{4}F(q^{-1})+x_{\text{L}}^{3}q^{4}b_{3}\right]$ (3)
is the reduced third virial coefficient of the binary mixture [74], with
$F(q)\equiv$ $(1+q)^{4}\cos^{-1}\frac{q^{2}+2q-1}{(1+q)^{2}}+8(1+q)^{2}\cos^{-1}\frac{1}{1+q}$
$-2\sqrt{q(2+q)}(3+2q+q^{2}).$ (4)
The parameters $\mu$ and $\lambda$ defined in Eqs. (2) are motivated by the “surplus” approximation for the equation of state of HD or HS fluid mixtures [75, 48, 74, 76].
According to Eqs. (2), $\phi_{\text{mixt}}$ depends linearly on $\mu$, while $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ depends linearly on $\lambda$. As discussed in the Appendix, these two assumptions again uniquely determine the structure of Eqs. (2). Figure 2 shows that these simple linear relationships are well satisfied with the standard value $\phi_{\text{mono}}=0.844$, particularly for $q=3$, and that the collapse of data from different mixtures is significantly improved compared with Brouwers’ model.
A complementary proposal for HD mixtures has recently been put forward by Zaccone [68, 69]. In this approach, the RCP fraction is determined by equating the compressibility factors of the mixture and the monodisperse system, $Z_{\text{mixt}}(\phi_{\text{mixt}})=Z_{\text{mono}}(\phi_{\text{mono}})$. Using the Scaled Particle Theory expressions [77, 78, 79], this leads to the explicit formula [69]:
$\phi_{\text{mixt}}=$ $1-\frac{(1-\phi_{\text{mono}})^{2}}{2}\Bigg[1-m_{2}^{-1}$
$+\sqrt{(1-m_{2}^{-1})^{2}+\frac{4m_{2}^{-1}}{(1-\phi_{\text{mono}})^{2}}}\Bigg].$ (5)
Figure 3 compares simulation data with our proposal, Brouwers’ original model, and Zaccone’s approach by plotting $\phi_{\text{mixt}}$ as a function of the area fraction of large disks,
$c_{\text{L}}=\frac{x_{\text{L}}q^{2}}{x_{\text{L}}q^{2}+x_{\text{S}}},$ (6)
for $q=1.4$, $1.7$, 22, and $3$. As observed, Brouwers’ proposal [Eq. (1a)] with the fitted value $\phi_{\text{mono}}=0.8425$ is accurate for $q=1.4$ and reasonably good for $q=1.7$, by construction, but fails for $q=2$ and $q=3$. Zaccone’s approach [Eq. (II)] with $\phi_{\text{mono}}=0.844$ performs fairly well for $q=1.4$, but underestimates $\phi_{\text{mixt}}$ for larger size ratios. Using the monodisperse value $\phi_{\text{mono}}\simeq 0.886$ [80] in Eq. (II), as done in Refs. [68, 69], leads to even poorer estimates. In contrast, our third-virial-based proposal provides the best agreement across all size ratios, demonstrating its effectiveness in capturing the universality of RCP in binary HD mixtures.
III Extension to polydisperse mixtures
Figure 4: Plot of the parameter $\mu$ as a function of the dispersity parameter $s$ for a polydisperse mixture with a log-normal size distribution.
We conjecture that Eq. (2a) remains a good candidate for estimating the RCP fraction of a polydisperse HD mixture characterized by a continuous size distribution $f(\sigma)$. In this case, the moments of the distribution are defined as $M_{n}=\int_{0}^{\infty}d\sigma,f(\sigma)\sigma^{n}$, and the reduced third virial coefficient is given by [74]
$\bar{B}_{3}=$ $\frac{8}{\pi M_{2}^{2}}\int_{0}^{\infty}d\sigma_{1}\int_{0}^{\infty}d\sigma_{2}\int_{0}^{\infty}d\sigma_{3}\,f(\sigma_{1})f(\sigma_{2})f(\sigma_{3})$
$\times\sigma_{12}^{2}\,\mathcal{S}_{\sigma_{13},\sigma_{23}}(\sigma_{12}),$ (7)
where
$\mathcal{S}_{a,b}(r)=$ $a^{2}\cos^{-1}\frac{r^{2}+a^{2}-b^{2}}{2ar}+b^{2}\cos^{-1}\frac{r^{2}+b^{2}-a^{2}}{2br}$
$-\frac{1}{2}\sqrt{2r^{2}(a^{2}+b^{2})-(b^{2}-a^{2})^{2}-r^{4}}.$ (8)
is the intersection area of two circles of radii $a$ and $b$ whose centers are separated by a distance $r$.
As an illustration, consider the log-normal distribution:
$f(\sigma)=\frac{1}{\sigma\sqrt{2\pi s^{2}}}e^{-[\ln(\sigma/\sigma_{0})]^{2}/2s^{2}},$ (9)
for which the moments are $M_{n}=\sigma_{0}^{n}e^{n^{2}s^{2}/2}$ and the reduced moments $m_{n}=e^{n(n-1)s^{2}/2}$. The value of $\bar{B}_{3}$ must then be evaluated numerically via the triple integral in Eq. (7). Figure 4 shows the parameter $\mu$ as a function of the dispersity parameter $s$. According to Eq. (2a), the RCP fraction $\phi_{\text{mixt}}$ is expected to vary approximately linearly with $\mu(s)$.
In general, the values of $\phi_{\text{mono}}$ and $\phi_{\text{mixt}}$ may depend on the packing protocol employed [27, 37]. Nevertheless, Eqs. (2) are expected to remain applicable provided that $\phi_{\text{mono}}$ and $\phi_{\text{mixt}}$ are determined using equivalent protocols.
IV Conclusions
In this work, we have proposed a simple yet effective modification of Brouwers’ geometric approach for estimating the RCP fraction of binary HD mixtures. By introducing a parameter based on the mixture’s reduced third virial coefficient, we obtained a nearly linear dependence of $\phi_{\text{mixt}}$ on $\mu$ and of $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ on $\lambda$, leading to a significantly improved collapse of simulation data across a wide range of size ratios. Comparisons with Brouwers’ original model and Zaccone’s proposal demonstrate that our approach provides the most accurate description of the RCP fraction for both small and moderate dispersity.
Furthermore, the method can be naturally extended to polydisperse mixtures with continuous size distributions, such as log-normal distributions, by computing the corresponding moments and third virial coefficient. The resulting parameter $\mu$ offers a universal measure of dispersity, allowing $\phi_{\text{mixt}}$ to be predicted as a nearly linear function of $\mu$.
Although this extension to polydisperse mixtures remains a conjecture, the excellent agreement with simulation results in the binary case suggests that it is worth testing. Its simplicity and predictive potential may stimulate new simulations to either confirm or refute the conjecture, providing a clear avenue for future research.
Overall, our results suggest that the third-virial-based parameter provides a robust and practically useful framework for understanding the universality of RCP in HD mixtures, with potential applications to more complex and polydisperse systems.
Acknowledgements.
A.S acknowledges financial support from Grant No. PID2024-156352NB-I00 funded by MCIU/AEI/10.13039/501100011033/FEDER, UE and from Grant No. GR24022 funded by the Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe.”
Data availability
Data supporting figures and numerical results are available from the corresponding author upon reasonable request.
Appendix
Here we show that Eqs. (1) are the only possible forms if one assumes that $1/\phi_{\text{mixt}}$ is a linear function of $\mu_{B}$ and that $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ is a linear function of $\lambda_{B}\equiv 1/(1-\mu_{B})$. Under these assumptions one may write
$\frac{1}{\phi_{\text{mixt}}}=\frac{1}{\phi_{\text{mono}}}-A\mu_{B},$ (10a)
$\frac{\phi_{\text{mixt}}}{1-\phi_{\text{mixt}}}=\frac{\phi_{\text{mono}}}{1-\phi_{\text{mono}}}+B\left(\lambda_{B}-1\right),$ (10b)
where we have used that $\mu_{B}=0$ in the monocomponent limit. Combining Eqs. (10) one obtains
$1-\left(\frac{1-\phi_{\text{mono}}}{\phi_{\text{mono}}}\right)^{2}\frac{B}{A}+\left(\frac{1-\phi_{\text{mono}}}{\phi_{\text{mono}}}B-1\right)\mu_{B}=0.$ (11)
Since this identity must hold for arbitrary $\mu_{B}$, both the constant and linear terms must vanish independently. This yields $A=1/B=(1-\phi_{\text{mono}})/\phi_{\text{mono}}$, in agreement with Eqs. (1).
An analogous reasoning applies to our proposal. If one assumes that $\phi_{\text{mixt}}$ is a linear function of $\mu$ and that $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ is a linear function of $\lambda\equiv 1/(1-\mu)$, one can write
$\phi_{\text{mixt}}=\phi_{\text{mono}}+C\mu,$ (12a)
$\frac{\phi_{\text{mixt}}}{1-\phi_{\text{mixt}}}=\frac{\phi_{\text{mono}}}{1-\phi_{\text{mono}}}+D\left(\lambda-1\right).$ (12b)
Eliminating $\phi_{\text{mixt}}$ between Eqs. (12) gives
$1-(1-\phi_{\text{mono}})^{2}\frac{D}{C}+\left[(1-\phi_{\text{mono}})D-1\right]\mu=0.$ (13)
Requiring this expression to be valid for all $\mu$ leads to $C=1/D=1-\phi_{\text{mono}}$, in agreement with Eqs. (2).